OpenStudy (anonymous):

Simplify 4j^2k^6(2j^11)^3k^5

OpenStudy (zzr0ck3r):

\[4j^2k^62j^{11^{(3k)^5}}\] this is what I read

OpenStudy (anonymous):

no sorry it is 4j^2k^6(2j^11)^3*k^5

OpenStudy (zzr0ck3r):

\[4j^22k^6(2j^{11})^3k^5?\]

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

\[4j^22k^6(2j^{11})^3k^5=4j^22k^6(2^3j^{33})k^5=64k^{5+6}j^{33+2}=64k^{11}j^{35}\]

OpenStudy (timaashorty):

well wow you are quick.

OpenStudy (zzr0ck3r):

lol

OpenStudy (timaashorty):

I was still working it out. lol

OpenStudy (zzr0ck3r):

well did you get the same thing?

OpenStudy (timaashorty):

Yup I agree

OpenStudy (anonymous):

Thank you for the answer and showing your work.